# Assumption-lean and Data-adaptive Post-Prediction Inference

### v3

@@ Abstract
A primary challenge facing modern scientific research is the limited availability of gold-standard data which can be both costly and labor-intensive to obtain. With the rapid development of machine learning (ML), scientists have relied on ML algorithms to predict these gold-standard outcomes with easily obtained covariates. However, these predicted outcomes are often used directly in subsequent statistical analyses, ignoring imprecision and heterogeneity introduced by the prediction procedure. This will likely result in false positive findings and invalid scientific conclusions. In this work, we introduce an assumption-lean and data-adaptive POst-Prediction Inference (POP-Inf) procedure that allows valid and powerful inference based on ML-predicted outcomes. Its “assumption-lean” property guarantees reliable statistical inference without assumptions on the ML-prediction, for a wide range of statistical quantities. Its “data-adaptive” feature guarantees an efficiency gain over existing post-prediction inference methods, regardless of the accuracy of ML-prediction. We demonstrate the superiority and applicability of our method through simulations and large-scale genomic data.
@@ 1 Introduction
A fundamental challenge in modern scientific research is the acquisition of gold standard data (Wang et al.,, 2023). These data, with their high accuracy and reliability, are essential to the validity of scientific discoveries, but obtaining them is often costly and labor-intensive. Fortunately, the advent and rapid development of machine learning (ML) has made it possible to predict outcomes using accessible covariates (LeCun et al.,, 2015; He et al.,, 2016). A prominent example is AlphaFold (Jumper et al.,, 2021), which uses readily available protein amino acid sequences to accurately predict protein structures that traditionally require extensive experimental efforts to determine. This ML-based approach has demonstrated its potential to substantially reduce the time and resources required to measure gold standard data (Stokes et al.,, 2020; Cheng et al.,, 2023).
Despite these benefits, replacing gold standard data with ML-prediction introduces new challenges, particularly in maintaining the validity of downstream statistical analyses. The indiscriminate use of such predictions, without acknowledging their distinction from observed gold-standard data, can lead to biased results and misleading scientific conclusions (Wang et al.,, 2020). This issue is exemplified by the statistical analysis using imputed gene expression in the Genotype-Tissue Expression (GTEx) project.
Example 1 (Multi-tissue Prediction of Gene Expression in GTEx Project).
Integrating gene expression across tissues and cell types is critical to understanding the coordinated biological mechanisms that drive disease and characterize homeostasis (Lonsdale et al.,, 2013). The GTEx project is a comprehensive public resource that studies gene expression and regulation in multiple human tissues (GTEx Consortium et al.,, 2015). It aims to improve our understanding of gene regulation and thus provide insights into the biological mechanisms behind various diseases and traits. However, many types of tissue are difficult to access and are not collected from each GTEx individual. The percentage of individuals with missing gene expression ranges from 5% to 90% (median 47%) across all tissues in GTEx (GTEx Consortium,, 2020). This leads to an incomplete picture of an individual’s physiological state and requires effective methods for multi-tissue prediction. To address this problem, several approaches have been proposed to impute gene expression in the uncollected tissue (Wang et al.,, 2016; Viñas et al.,, 2023; Basu et al.,, 2021). However, after prediction, imputed gene expression has been treated directly as observed gene expression and used in subsequent statistical analysis for scientific discovery (Viñas et al.,, 2023), such as understanding the genetic basis of gene regulation. As demonstrated in (Wang et al.,, 2020), direct use of the estimation results is likely to produce false positives, leading to invalid scientific conclusions.
The problem of making inferences after acquiring ML predictions is not limited to the specific field in Example 1, but is instead a pervasive problem in various domains where ML is applied (Bullock et al.,, 2020). The inherent complexity and “black box” nature of many ML models adds layers of difficulty in validating the statistical inferences derived after prediction. This highlights the need for methods that can effectively bridge the gap between ML-prediction and the demands of scientifc rigor, ensuring that the integration of ML into research does not compromise the validity of scientific conclusions.
@@
To address this problem, we introduce a simple, valid, powerful, and widely applicable approach called Post-Prediction Inference (POP-Inf) that is assumption-lean and data-adaptive. POP-Inf is designed to integrate ML-prediction with the observed gold standard data for a general estimation problem. We highlight two key features of our POP-Inf method:
@@
  • Assumption-lean: The POP-Inf estimator is consistent and asymptotically normal, without making any assumptions about the ML prediction algorithm. This means that POP-Inf can be used with arbitrarily misspecified ML-prediction.
@@
  • Data-adaptive: The POP-Inf estimator is adaptive to the accuracy of the ML-prediction. It utilizes more information from a “good” ML-prediction to reduce variance, and avoids variance inflation otherwise. It degenerates to many existing estimators under certain cases of the ML-prediction.
@@
Our contributions are threefold: (i) We present a simple and data-adaptive method of post-prediction inference for an estimation problem defined through estimating equations, without any assumptions on the ML predictions or any parameterization of the true data generating process. (ii) We establish the consistent and asymptotic normality of our proposed estimator. We further demonstrate the optimality of our estimator in terms of asymptotic variance over a class of estimators that includes many existing estimators in the post-prediction inference. (iii) We connect our method with the theory of semiparametric efficiency. We derive the semiparametric efficiency bound and demonstrate the potential of our proposed method in achieving that bound. In addition, once the nuisance function cannot be correctly specified, an efficient influence function-guided estimator with estimated nuisances plugged in can be less efficient than the proposed POP-Inf procedure.
@@
The rest of this paper is organized as follows. In Section 2, we formulate our problem and illustrate our method with mean estimation to build an intuition. Next in Section 3, we introduce a general protocol for applying POP-Inf to estimands defined through estimating equations. Further, we establish the asymptotic properties of the POP-Inf estimator and describe the estimation procedure. We build the connection with the semiparametric efficiency theory in Section 4 and describe extensions in Section 5. Through extensive simulations (Section 6) and real data analysis (Section 7), we validate our theoretical claims and demonstrate the practical utility and effectiveness of the POP-Inf estimator. This paper concludes with a discussion in Section 8.
@@ 2 Problem Formulation
Let $Y$ be a scalar outcome variable and ${\bf X}$ be a $d$-dimensional covariate variable. The scientific interest is to estimate a $q$-dimensional parameter ${\bm{\theta}}$ defined through an estimating equation,
@@
where $\bm{\psi}(\cdot,\cdot;{\bm{\theta}})$ is a user-defined function. Such a definition of ${\bm{\theta}}$ is very general including outcome mean, outcome quantile, least squares coefficients, or any other specific quantities of interest involving both ${\bf x}$ and $y$, including the unique minimizer of a loss function and the maximizer of a criterion function (Van der Vaart,, 2000) ; see Section 2.1 for an example of estimating the outcome mean.
@@
We consider the availability of an external and independent prediction algorithm $\widehat{f}(\cdot)$ on ${\bf z}$ that produces predictions $\widehat{f}({\bf z})$ for the outcome variable $Y$. We assume that the operating characteristics of $\widehat{f}$ are unknown to the end-user, and the data used to fit $\widehat{f}$ are unavailable. Thus $\widehat{f}$ is considered a black box function. In addition, this $\widehat{f}$ is can be incorrect and even biased for predicting $y$.
With access to this black-box $\widehat{f}$, our primary goal is to improve efficiency while making inference on the parameter ${\bm{\theta}}$. We refer to our problem as post-prediction inference in that the inference is made after predictions having been made independently (Wang et al.,, 2020). Unlike existing work that provides tools to estimate the true “$f$” or assuming it can be well approximated, we make no assumptions on $\widehat{f}$. The resulting protocol can benefit researches where training a prediction algorithm is unrealistic (e.g., training ChatGPT), and provide guidelines for scientists who are not equipped with computational resources or knowledge about building prediction algorithms.
@@
We deliberately split the covariate into two (possibly overlapping) parts, ${\bf Z}$ and ${\bf X}$, where the former is predictive of the outcome $Y$ and the latter appears in the estimating equation (1). They could be the same, but a wide range of applications suggest they are different. When ${\bf Z}$ is not a subset of ${\bf X}$, the information it carries about the outcome $Y$ may not be informative for the estimation problem (1). Hence, a better prediction algorithm $\widehat{f}$ does not necessarily correspond to a better estimate of ${\bm{\theta}}$. As justified above, this work does not attempt to improve the prediction algorithm, but to empower the inference on the parameter of interest, given any arbitrary black-box $\widehat{f}$.
In the followings, Section 2.1 illustrates the key idea of the proposed method by an example of estimating outcome mean. Section 2.2 reviews related work. After that, in Section 3, we introduce the proposed framework for the estimation problem defined in equation (1).
@@ 2.1 Example: estimating the outcome mean
We provide the intuition of our approach using mean estimation. Consider $\psi(y,{\bf x};{\bm{\theta}})=y-\theta$. The resulting estimand is $\theta=\mathbb{E}(Y)$, the outcome mean. For illustration propose, we consider ${\bf z}={\bf x}$. Then the classic approach calculates an estimator from the sample average of observed outcomes,
@@
This estimator is unbiased, but can have high variance due to the small size of labeled data. To increase efficiency with the ML-predicted outcomes in unlabeled data, we propose to introduce another unbiased estimator of zero with a weight scalar $\omega$,
@@
$\widehat{\theta}_{\textnormal{POP-Inf}}(\omega)=\dfrac{1}{n}\sum_{i=1}^{n}y_{i}+\omega\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}\widehat{f}({\bf z}_{i})-\dfrac{1}{n}\sum_{i=1}^{n}\widehat{f}({\bf z}_{i})\right]$ (3)
We refer to this estimator as the POP-Inf estimator.
@@
Both the classic estimator and recently-introduced prediction-powered (PP) estimator (Angelopoulos et al., 2023a, ) can be viewed as special cases of the POP-Inf. The classic estimator is the POP-Inf estimator with $\omega=0$. The PP estimator is the POP-Inf estimator with $\omega=1$:
$\widehat{\theta}_{\textnormal{PP}}=\dfrac{1}{n}\sum_{i=1}^{n}y_{i}+\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}\widehat{f}({\bf z}_{i})-\dfrac{1}{n}\sum_{i=1}^{n}\widehat{f}({\bf z}_{i})\right].$ (4)
@@
As detailed in Section 4, another candidate estimator is guided from the efficient influence function (EIF). We refer to this estimator as $*{}^{*}$-based where $*{}^{*}$ represents the fact that the nuisance function is plugged in as an estimate,
$\widehat{\theta}_{\textnormal{EIF}^{*}}=\dfrac{1}{n}\sum_{i=1}^{n}y_{i}+\dfrac{N}{N+n}\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}\widehat{f}({\bf z}_{i})-\dfrac{1}{n}\sum_{i=1}^{n}\widehat{f}({\bf z}_{i})\right].$ (5)
@@
It is also a special case of the POP-Inf estimator with $\omega=N/(n+N)$. This estimator achieves the semiparametric efficiency bound when $\widehat{f}$ approximates its true counterpart at a sufficiently fast rate.
@@
Since all these estimators are summations of an unbiased estimator of the population mean and an unbiased estimator of zero, they are consistent for the population mean. In addition, Wald-type confidence intervals can be constructed using the following asymptotic variance,
@@
$\underbrace{\frac{\textnormal{Var}_{l}(Y)}{n}+\underbrace{\left(\omega^{2}\frac{\textnormal{Var}_{l}[\widehat{f}]}{n}-2\omega\frac{\textnormal{Cov}_{l}[Y,\widehat{f}]}{n}+\omega^{2}\frac{\textnormal{Var}_{u}[\widehat{f}]}{N}\right)}_{\text{Additional terms}}}_{\text{Variance of {POP-Inf}\leavevmode\nobreak\ estimator}},$ (6)
where $\textnormal{Var}_{l}[\cdot]$ and $\textnormal{Cov}_{l}[\cdot,\cdot]$ represent the variance and covariance on the labeled, respectively, and $\textnormal{Var}_{u}[\cdot]$ is the variance on the unlabeled. Note that although $\textnormal{Var}_{l}[\cdot]$ and $\textnormal{Var}_{u}[\cdot]$ are essentially the same, we still use different notations to indicate their different sample analogs that will guide the practice.
@@
To guarantee that the asymptotic variance (6) is no larger than the classic approach regardless of the quality of $\widehat{f}$, the additional terms need to be no larger than zero. A key insight is that these additional terms are a quadratic function of $\omega$,
$q(\omega)=\omega^{2}\left\{\frac{\textnormal{Var}_{l}[\widehat{f}]}{n}+\frac{\textnormal{Var}_{u}[\widehat{f}]}{N}\right\}-2\omega\frac{\textnormal{Cov}_{l}[Y,\widehat{f}]}{n},$ (7)
@@
which achieves its minimum at $\omega_{0}=\textnormal{Cov}_{l}[Y,\widehat{f}({\bf z})]/\{\textnormal{Var}_{l}[\widehat{f}]+n\textnormal{Var}_{u}[\widehat{f}]/N\}$ with minimum value
$q(\omega_{0})=\dfrac{-\textnormal{Cov}_{l}[Y,\widehat{f}]}{n\textnormal{Var}_{l}[\widehat{f}]+n^{2}\textnormal{Var}_{u}[\widehat{f}]/N}\leq 0,$ (8)
@@
where the equality holds if and only if $\textnormal{Cov}_{l}[Y,\widehat{f}]=0$. In this case, our POP-Inf estimator is no less efficient than the classic estimator. In contrast, both PP and $*{}^{*}$-based estimator may be less efficient than the classic estimator, when the ML predictions are not accurate. As we shall see in Section 3, similar patterns appear in estimating the general parameter defined in (1). In addition, the additional terms are below zero if and only if $\textnormal{Cov}_{l}[Y,\widehat{f}]>\omega(\textnormal{Var}_{l}[\widehat{f}]+n\textnormal{Var}_{u}[\widehat{f}]/N)/2$. This indicates that for efficiency improvement, the prediction accuracy of the ML algorithm cannot be low.
@@
We provide an intuition for the “data-adaptive” feature of POP-Inf. When the ML algorithm is purely random, $\textnormal{Cov}_{l}[Y,\widehat{f}]=0$ and hence the optimal weight $\omega_{0}=0$. Therefore, POP-Inf degenerates to the classic estimator using labeled data only. Such choice suits the expectation that one should not use this ML-prediction. On the other hand, if the ML-prediction is perfect in that the predictions are exactly the true outcomes, $\textnormal{Cov}_{l}[Y,\widehat{f}]=\textnormal{Var}_{l}[\widehat{f}]=\textnormal{Var}_{l}[Y]$ and hence the optimal weight is $\omega_{0}=N/(N+n)$. In this case, POP-Inf degenerates to the EIF-based estimator that achieves the semiparametric efficiency bound. This is also intuitive since perfect ML-prediction can be treated as gold-standard data and the estimator should utilize all the (measured and predicted) information and thus is efficient.
Figure 1 displays numeric results for estimating outcome mean and ordinary least squares coefficients with details contained in Section 6. With a common $\widehat{f}$ obtained from an independent sample, we construct the POP-Inf estimator, the $*{}^{*}$-based estimator, the PP estimator, and the classic estimator. On the x-axis, each $\gamma$ corresponds to a data generating system which roughly speaking quantifies the capability of $\widehat{f}$ for predicting $y$: the higher $\gamma$ is, the better $\widehat{f}$ would be. On the y-axis, left panels display the coverage rates; right panels coordinate the ratio of averaged confidence interval widths across $1000$ simulations between the classic estimator and one of the competing estimators.
All methods yield satisfying coverage rates centered around the nominal level $95\%$, but their efficiency can be quite different. From right panels, the $*{}^{*}$-based estimator and the PP estimator perform comparably, both less efficient than the classic approach when $\gamma$ is small, and more efficient as $\gamma$ becomes close to $0.5$. On the other hand, the POP-Inf estimator is always equally or more efficient than the classic approach despite the value of $\gamma$. In addition, the POP-Inf estimator is more efficient than the $*{}^{*}$-based and PP estimators across $\gamma\in[0,0.5]$. Additional results in Section 6 corroborate such patterns. From empirical evidence, the POP-Inf estimator is desirable in practice.
Figure 1: Relative efficiency and coverage rates in simulations with $\gamma\in[0,0.5]$, $n=250$ and $N=20n$. In each panel, the x-axis coordinates the data generating parameter $\gamma$ detailed in Section 6. Left panels plot the ratios between averaged standard errors of a competing method and the classic approach, with dashed horizontal line intersects the y-axis at $1$. A ratio lower than one represents a smaller standard error than the classic approach. Right panels plot coverage rates with dashed horizontal line intersects the y-axis at the nominal level $0.95$.

### v4

@@ Abstract
A primary challenge facing modern scientific research is the limited availability of gold-standard data which can be costly, labor-intensive, or invasive to obtain. With the rapid development of machine learning (ML), scientists can now employ ML algorithms to predict gold-standard outcomes with variables that are easier to obtain. However, these predicted outcomes are often used directly in subsequent statistical analyses, ignoring imprecision and heterogeneity introduced by the prediction procedure. This will likely result in false positive findings and invalid scientific conclusions. In this work, we introduce PoSt-Prediction Adaptive inference (PSPA) that allows valid and powerful inference based on ML-predicted data. Its “assumption-lean” property guarantees reliable statistical inference without assumptions on the ML prediction. Its “data-adaptive” feature guarantees an efficiency gain over existing methods, regardless of the accuracy of ML prediction. We demonstrate the statistical superiority and broad applicability of our method through simulations and real-data applications.
@@ 1 Introduction
Gene expression data obtained from various tissue and cell types can provide crucial insights into the coordinated biological mechanisms that drive disease etiology and characterize homeostasis (Lonsdale et al.,, 2013). However, some important tissues are often difficult to collect, which leads to underwhelming size of gene expression samples in those tissues. For example, the Genotype-Tissue Expression (GTEx) project is a comprehensive study focusing on gene expression regulation in many human tissues (GTEx Consortium et al.,, 2015). The percentage of individuals with missing gene expression ranges from 5% to 90% (median 47%) across all tissues in GTEx (GTEx Consortium,, 2020). This limits the scientific understanding of transcription regulation across tissue contexts.
The difficulty in obtaining gold-standard data certainly extends beyond human gene expression applications (Wang et al.,, 2023). While gold-standard data with high reliability are essential to the validity of scientific discoveries, obtaining them is often costly and labor-intensive. Fortunately, the advent and rapid development of machine learning (ML) have enabled prediction of outcomes using more accessible variables (LeCun et al.,, 2015), showing great potential in reducing the need to acquire gold-standard data.
Despite these benefits, replacing gold-standard data with ML predictions introduces new challenges, particularly in maintaining the validity of downstream statistical analyses. This issue is exemplified by the statistical analysis using imputed gene expression in the GTEx project. To address the insufficient sample size for rare tissues, several approaches have been proposed to impute gene expressions in these tissues using expression data from tissues that are easier to acquire (Wang et al.,, 2016; Viñas et al.,, 2023; Basu et al.,, 2021). However, imputed gene expression is often treated as if it were observed, and used directly in subsequent statistical analyses for scientific discovery (Viñas et al.,, 2023), such as exploring sex differences and the genetic basis of gene regulation. As demonstrated in Wang et al., (2020); Angelopoulos et al., 2023a , direct use of ML prediction in statistical inference will most likely produce false positives findings, leading to invalid scientific conclusions.
The challenge of making valid inferences from ML predictions extends beyond the specific example discussed, affecting numerous scientific disciplines where ML is applied (Bullock et al.,, 2020). Recent work has introduced a method called prediction-powered inference (PPI), wwhich leverages a small set of labeled data with gold-standard outcomes and a large amount of unlabeled data with ML predictions to perform valid statistical inference (Angelopoulos et al., 2023a, ). Although validity is guaranteed, PPI may be less efficient than the inference relying solely on labeled data. For instance, when applied to GTEx data, PPI  identifies fewer sex-biased gene expressions compared to the classical approach based on labeled data alone, highlighting a loss of statistical efficiency (Figure 1). This limitation affects the broader applicability of PPI and similar methods.
@@
To address this problem, we introduce a valid, powerful, and widely applicable approach named PoSt-Prediction Adaptive inference (PSPA). PSPA is designed to integrate ML prediction with the observed gold-standard data to ensure valid and efficient statistical inference. We highlight two key features of our PSPA method:
@@
  • Assumption-Lean: The PSPA estimator is consistent and asymptotically normal, ensuring valid inference with no assumptions on the ML model and prediction accuracy. This means that PSPA can be used with arbitrarily misspecified ML prediction.
@@
  • Data-Adaptive: The PSPA estimator is adaptive to the accuracy of the ML prediction. It utilizes more information from a “good” ML prediction to reduce variance and avoids variance inflation otherwise. Below, we will demonstrate that PSPA achieves element-wise variance reduction compared to existing methods, regardless of the ML prediction quality.
@@
Our contributions are threefold: (i) We present a simple and data-adaptive method of post-prediction inference for an estimation problem defined through estimating equations, without any assumptions on the ML prediction or any parameterization of the true data-generating process. (ii) We establish the consistent and asymptotic normality of our proposed estimator. We further demonstrate the optimality of our estimator in terms of asymptotic variance over a class of estimators that includes existing estimators in post-prediction inference. (iii) In contrast to current methods which only focus on incorporating ML-predicted labels in statistical inference, our approach extends to utilizing both ML-predicted covariates and labels, enhancing the versatility and applicability of post-prediction inference.
@@
The rest of this paper is organized as follows. In Section 2, we formulate our problem and illustrate our method with mean estimation to build an intuition. Next in Section 3, we introduce a general protocol for applying PSPA to estimands defined through estimating equations. Further, we establish the asymptotic properties of the PSPA estimator and describe the estimation procedure. We build the connection with the semiparametric efficiency theory in Section 4 and describe extensions in Section 5. Through extensive simulations (Section 6) and real data analysis (Section 7), we validate our theoretical claims and demonstrate the practical utility and effectiveness of the PSPA estimator. We then compared our method with a recent method PPI++ (Angelopoulos et al., 2023b, ) in Section 8. This paper concludes with a discussion in Section 9.
Figure 1: Comparison of PPI, PSPA, and classical method in identifying sex-biased gene expressions using GTEx data. (a) number of sex-biased genes identified by each of the four approaches. (b) x-axis: absolute value of imputation correlation. y-axis: relative ratio of estimated standard error between PPI and classical method. (c) same as b but for our method PSPA. The dashed line represents $y=1$ in (b)-(c).
@@ 2 Problem Formulation
Let $Y$ be a scalar outcome and ${\bf X}$ be $d$-dimensional covariates. The scientific interest is to estimate a $q$-dimensional parameter ${\bm{\theta}}$ defined through an estimating equation,
@@
where $\bm{\psi}(\cdot,\cdot;{\bm{\theta}})$ is a user-defined function. Such a definition of ${\bm{\theta}}$ is very general including outcome mean, outcome quantile, least squares coefficients, or any other specific quantities of interest involving both ${\bf x}$ and $y$, including the unique minimizer of a loss function and the maximizer of a criterion function (Van der Vaart,, 2000).
@@
We consider the availability of an external and independent prediction algorithm $\widehat{f}(\cdot)$ on ${\bf z}$ that produces predictions $\widehat{f}$ for the outcome variable $Y$. We assume that the operating characteristics of $\widehat{f}$ are unknown to the user, and the data used to fit $\widehat{f}$ are unavailable. $\widehat{f}$ is considered a black box function and can be incorrect in predicting $y$.
With access to this black-box $\widehat{f}$, our primary goal is to improve efficiency while making inference on the parameter ${\bm{\theta}}$. We refer to our problem as post-prediction inference, where inference occurs after predictions are made independently (Wang et al.,, 2020; Angelopoulos et al., 2023a, ). Specially, we make no assumptions on $\widehat{f}$. The resulting protocol can benefit researches where training a prediction algorithm is unrealistic, and offers guidelines for scientists lacking computational resources or predictive modeling expertise.
Example 1 (Sex-differentiated gene expressions).
We present the scientific problem that motivates our setup: identifying sex-biased gene expressions in the brain cortex tissue using data from the GTEx project. For this analysis, we have labeled data denoted by =${\cal L}=\{(y_{i},\widehat{f}_{i},{\bf x}_{i},{\bf z}_{i}),i=1,\cdots,205\}$, where $y_{i}$ represents the gene expression in brain cortex tissue, ${\bf x}_{i}$ includes a binary indicator for biological sex $in\{0,1\}$, along with other technical factors such as surrogate variables for batch effects, age, RNA integrity number, and total ischemic time. and total ischemic time. ${\bf z}_{i}$ is the whole-transcriptome gene expression profile in whole blood tissue that is easier to access and $\widehat{f}$ is the imputed gene expression for in the brain cortex tissue using ${\bf z}_{i}$. In addition, our unlabeled data, ${\cal U}=\{(\widehat{f}_{i},{\bf x}_{i},{\bf z}_{i}),i={206},\cdots,670\}$, contains 465 samples whose gene expression was measured in whole blood but not in brain cortex. Our interest lies in the linear regression coefficient ${\bm{\theta}}$ as a solution to $\mathbb{E}[{\bf x}(y-{\bf x}^{\rm T}{\bm{\theta}})]=0$, focusing particularly on the coefficients corresponding to biological sex.
@@
We split the covariate into two (possibly overlapping) parts, ${\bf Z}$ and ${\bf X}$, where the former is predictive of the outcome $Y$ and the latter appears in the estimating equation (1). They could be the same, but a wide range of applications suggest they are different. When ${\bf Z}$ is not a subset of ${\bf X}$, the information it carries about the outcome $Y$ may be the most informative for the estimation problem (1). Hence, a better prediction algorithm $\widehat{f}$ does not necessarily correspond to a better estimate of ${\bm{\theta}}$. As justified above, this work does not attempt to improve the prediction algorithm, but to empower the inference on the parameter of interest, given any arbitrary black-box $\widehat{f}$.
In the following, Section 2.1 presents the key idea of the proposed method using mean estimation as an example. Section 2.2 reviews related work. In Section 3, we introduce the proposed framework for the estimation problem defined in equation (1).
@@ 2.1 Example: estimating the outcome mean
We provide the intuition of our approach using mean estimation. Consider $\psi(y,{\bf x};{\bm{\theta}})=y-\theta$. The resulting estimand is $\theta=\mathbb{E}(Y)$, the outcome mean. Then the classical approach calculates an estimator from the sample average of observed outcomes,
@@
This estimator is unbiased, but can have high variance due to the small sample size of labeled data. To increase efficiency with the ML-predicted outcomes in unlabeled data, we propose to introduce another unbiased estimator of zero with a weighting scalar $\omega$,
@@
$\widehat{\theta}_{\textnormal{{{PSPA}}}}(\omega)=\dfrac{1}{n}\sum_{i=1}^{n}y_{i}+\omega\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}\widehat{f}-\dfrac{1}{n}\sum_{i=1}^{n}\widehat{f}\right]$ (3)
@@
We refer to this estimator as the PSPA estimator. Both the classical and PPI estimator (Angelopoulos et al., 2023a, ) can be viewed as special cases of the PSPA. The classical estimator is the PSPA estimator with $\omega=0$. The PPI estimator is the PSPA estimator with $\omega=1$:
$\widehat{\theta}_{\textnormal{PP}}=\dfrac{1}{n}\sum_{i=1}^{n}y_{i}+\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}\widehat{f}-\dfrac{1}{n}\sum_{i=1}^{n}\widehat{f}\right].$ (4)
@@
As detailed in Section 4, another candidate estimator is guided by the efficient influence function (EIF). We refer to this estimator as EIF-based where represents the fact that the nuisance function is plugged in as an estimate,
$\widehat{\theta}_{\textnormal{EIF}^{*}}=\dfrac{1}{n}\sum_{i=1}^{n}y_{i}+\dfrac{N}{N+n}\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}\widehat{f}-\dfrac{1}{n}\sum_{i=1}^{n}\widehat{f}\right].$ (5)
@@
This is a special case of PSPA estimator with $\omega=\frac{N}{N+n}$, achieving the semiparametric efficiency bound when $\widehat{f}$ approximates its true counterpart at a sufficiently fast rate.
@@
Since these estimators are summations of consistent estimators of the population mean and zero, they are consistent for the population mean. Additionally, Wald-type confidence intervals can be constructed using a consistent estimator for the asymptotic variance.
@@
$\underbrace{\frac{\operatorname{Var}(Y)}{n}+\underbrace{\left(\omega^{2}\frac{\operatorname{Var}[\widehat{f}]}{n}-2\omega\frac{\operatorname{Cov}[Y,\widehat{f}]}{n}+\omega^{2}\frac{\operatorname{Var}[\widehat{f}]}{N}\right)}_{\text{Additional terms}}}_{\text{Variance of {{PSPA}}\leavevmode\nobreak\ estimator}},$ (6)
@@
To guarantee that the asymptotic variance (6) is no larger than the classical approach regardless of the quality of $\widehat{f}$, the additional terms need to be no larger than zero. A key insight is that these additional terms are a quadratic function of $\omega$:
$q(\omega)=\omega^{2}\left\{\frac{\operatorname{Var}[\widehat{f}]}{n}+\frac{\operatorname{Var}[\widehat{f}]}{N}\right\}-2\omega\frac{\operatorname{Cov}[Y,\widehat{f}]}{n},$ (7)
@@
which achieves its minimum at $\omega^{\textnormal{opt}}=\operatorname{Cov}[Y,\widehat{f}]/\{\operatorname{Var}[\widehat{f}]+n\operatorname{Var}[\widehat{f}]/N\}$ with minima
$q(\omega^{\textnormal{opt}})=\dfrac{-\operatorname{Cov}[Y,\widehat{f}]}{n\operatorname{Var}[\widehat{f}]+n^{2}\operatorname{Var}[\widehat{f}]/N}\leq 0,$ (8)
@@
where the equality holds if and only if $\operatorname{Cov}[Y,\widehat{f}]=0$. In this case, our PSPA estimator is no less efficient than the classical estimator. In contrast, both PPI and EIF-based estimators can be less efficient than the classical estimator, when the ML predictions are not accurate. As we shall see in Section 3, similar patterns appear in estimating the general parameter defined in (1). In addition, the additional terms are below zero if and only if $\operatorname{Cov}[Y,\widehat{f}]>\frac{\omega(N+n)\operatorname{Var}[\widehat{f}]}{2N}$. This indicates that in order for PPI to achieve efficiency improvement (where $\omega=1$), the prediction accuracy of the ML algorithm cannot be low.
@@
We provide an intuition for the “data-adaptive” feature of PSPA. When the ML algorithm is purely random, $\operatorname{Cov}[Y,\widehat{f}]=0$ and hence the optimal weight $\omega^{\textnormal{opt}}=0$. Therefore, PSPA degenerates to the classical estimator that uses only labeled data. This suits the expectation that such ML prediction should not be used in inference. On the other hand, if the ML prediction is perfect, $\operatorname{Cov}[Y,\widehat{f}]=\operatorname{Var}[\widehat{f}]=\operatorname{Var}[Y]$ and hence the optimal weight is $\omega^{\textnormal{opt}}=N/(N+n)$. In this case, PSPA degenerates to the EIF*-based estimator that achieves the semiparametric efficiency bound. This is also intuitive since perfect ML prediction can be treated as gold-standard data and the estimator should utilize all the (measured and predicted) information and thus is efficient.